i(iota)to the power i(iota).

Grade:12

3 Answers

vikas askiitian expert
13 years ago

z = ei@ (representation of complex number Z in ellulers form)

z1 = ii (given)

taking log both sides

logz1 = ilogi  .................1

now , logi = log(ei@)

= log(eipi/2)

= ipi/2

puting this value of logi in eq 1

logZ1 = i2(pi/2) = -pi/2  (i2 = -1)

Z1 = e(-pi/2) , this is the required result

Tejas Paresh Lodaya
12 years ago

Hehe log is only defined for non negative real no. :P :D
And you solved it? 

Vikas soni
5 years ago
i= e^i(π÷2)Now,i^i= {e^(i(π÷2))}^i= e^(-(π÷2)) which is positive and real. So the upper provided solution is true